OpenStudy (anonymous):

Create an expression to find the time for the given distance and rate, and find the value of the expression.

OpenStudy (anonymous):

Create an expression to find the time for the given distance and rate, and find the value of the expression.

OpenStudy (anonymous):

d=84 feet; r= 12 feet/second

OpenStudy (anonymous):

thats the eqation

OpenStudy (anonymous):

help please

OpenStudy (linyu):

\[v=\frac{ dx }{ dt }\] \[dx=vdt\] \[\int\limits\limits_{x _{0}}^{x}dx=\int\limits\limits_{0}^{\Delta t}{ vdt }\] \[\int\limits_{x _{0}}^{x} x=\int\limits_{0}^{\Delta t}(v _{0}+at)dt\] \[x-x _{0}=v _{0} \Delta t+\frac{ 1 }{ 2 }at ^{2}\] since\[a=0, x _{0}=0,v _{0}=r,x=d,\Delta t=t\] therefore\[d=rt\]\[t=\frac{ d }{ r }\]

OpenStudy (anonymous):

woah thanks

OpenStudy (linyu):

it was supposed to be 1/2a(Delta t)square sorry about that I wrote 1/2atsquare

OpenStudy (anonymous):

its alright